# A Unified Approach to Submodular Maximization Under Noise

### v1

@@ 4.2 Unconstrained Submodular Maximization
Finally, we consider maximizing any (non-monotone) submodular function without constraints. Here, we instantiate Theorem 3.4 with the double greedy algorithm [3] which is known to give a $1/2$-approximation for unconstrained submodular maximization. At a high-level, the double greedy algorithm works as follows. We first initialize two sets $X_{0}=\emptyset$ and $Y_{0}=N$. Let $N=\{e_{1},\ldots,e_{n}\}$. For $t=1,\ldots,n$, we check the marginal value $a_{t}$ of adding $e_{t}$ to $X_{t-1}$ and the marginal value $b_{t}$ of removing $e_{t}$ from $X_{t-1}$. With probability $a_{t}/(a_{t}+b_{t})$ (with some clipping operations if necessary), we add $e_{t}$ to $X_{t-1}$ to get $X_{t}$ and set $Y_{t}=Y_{t-1}$. Otherwise, we set $X_{t}=X_{t-1}$ and remove $e_{t}$ from $Y_{t-1}$ to get $Y_{t}$. A formal description of the algorithm can be found in Appendix D. The key fact that we require is the following lemma, with proof given in Appendix D.
@@
Fix $\varepsilon\in(0,\frac{2\nu^{2}}{\alpha})$. Suppose $n\geq\Omega(\frac{1}{\varepsilon}\log^{2}(\frac{n}{\varepsilon}))$. By letting the $\mathcal{A}$ in Algorithm 1 be the double greedy algorithm [3], we obtain a polynomial-time algorithm for maximizing unconstrained non-monotone submodular functions with noisy oracle satisfying $\mathbb{E}[f(\operatorname{ALG})]~\geq~\big(\frac{1}{2}-O(\varepsilon)\big)f(O^{*})$.
@@ 5 Simulation Results
  • Double greedy (DG) with exact value oracle [3]: this is a worst-case optimal polynomial-time algorithm in the noiseless setting. It uses the exact value oracle and is used only for reference.
@@ Appendix B Missing Proofs from Section 3
Lemma B.1 (Hoeffding’s inequality for sub-exponential distributions: see, e.g., page 29 in [24]).
@@
Fix a set $S\subseteq N\setminus H$. Let $\mu=F^{H,t}(S)=\mathbb{E}_{H^{\prime}\sim H[t]}f(S\cup H^{\prime})=\frac{1}{\binom{h}{t}}\sum_{H^{\prime}\in H[t]}f(S\cup H^{\prime})$. By definition, $H_{1},\ldots,H_{m}$ are sampled from $H[t]$ without replacement. Since each $f(S\cup H^{\prime})$ is bounded in $0,f_{\max}]$, by Hoeffding’s inequality for sampling without replacement (see, e.g., Proposition 1.2 in [1]), we have
@@
11Use pipage rounding [4] (which does not require access to $\hat{f}$) to convert the fractional solution $x(1)$ to a discrete set $S\in\mathcal{I}$.
@@
Algorithm 2 Measured continuous greedy [9] with approximate value oracle
In this section, we establish the robustness of the measured continuous greedy algorithm [9] (the full algorithm is given in Algorithm 2). Let $\hat{f}$ be an $\varepsilon$-approximate value oracle for $f$. The proof for the approximation ratios of the measured continuous greedy algorithm with exact value oracle is provided by Feldman [8]. We analyze how the $\varepsilon$-approximate value oracle $\hat{f}$ will affect the approximation ratios. The main step is Lemma C.4 which provides a replacement of Corollary 3.2.7 of [8]. Once this is established, we can apply the remaining arguments in Section 3.2.1 and Section 3.2.2 of [8] by replacing the discretization error of $O(n^{3}\delta^{2})$ with the discretization and approximation error established in Lemma C.4.
@@ C.2 The Non-Monotone Case
The following lemma can be established from following the proofs of Lemma 3.2.8, Lemma 3.2.9, Corollary 3.2.10, and Lemma 3.2.11 from [8] verbatim but replacing $O(n^{3}\delta)$ in their argument with $O(\alpha n+n^{3}\delta)$ (i.e. Lemma C.4).
@@ 7   With probability $\hat{p}_{i}$, let $X_{i}=X_{i-1}\cup\{u_{i}\},~Y_{i}=Y_{i-1}$;
Algorithm 3 Double greedy [3] with approximate value oracle
This section proves the robustness of the double greedy algorithm [3] (given in Algorithm 3) against approximate value oracle.
@@
In this case, it can be easily verified that $r_{i}=\max\{p_{i}b_{i},q_{i}a_{i}\}-\tfrac{1}{2}(p_{i}a_{i}+q_{i}b_{i})\leq 0$. Due to approximate oracle, $r_{i}=\max\{\hat{p}_{i}b_{i},\hat{q}_{i}a_{i}\}-\tfrac{1}{2}(\hat{p}_{i}a_{i}+\hat{q}_{i}b_{i})$ may be positive. [19] show that the loss of performance of the double greedy algorithm due to approximate oracle can be upper bounded by $\sum_{i=1}^{n}r_{i}$:
Lemma D.2 (see Theorem 2.1 in [19] or Lemma 4.3 in [11]).
@@
    On the one hand, $b\leq\hat{b}+\varepsilon_{b}\leq\varepsilon_{b}$. On the other hand, because $a+b\geq 0$ holds for any submodular function [3], we have $a\geq-b\geq-\varepsilon_{b}$. Therefore,
@@
    On the one hand, $a\leq\hat{a}+\varepsilon_{a}\leq\varepsilon_{a}$. On the other hand, because $a+b\geq 0$ holds for any submodular function [3], we have $b\geq-a\geq-\varepsilon_{a}$. Therefore,
@@ Appendix F Discussion on High Probability Results
For monotone submodular functions, Huang et al. [16] provide a method to do noisy comparison. For any set $S\subseteq N$, define the following comparison surrogate function $f_{0}(S)$ and the noisy version $\tilde{f}_{0}(S)$:
@@
For non-monotone submodular functions, the above approach to obtaining high-probability result does not work because: (1) the $f(S_{i})\geq f_{0}(S_{i})$ step in the noisy comparison analysis no longer holds, and (2) the size of the solution $|S_{i}|$ can be less than the rank $r$ of the matroid, so the $(1-\frac{1}{|S_{i}|})$ factor in the approximation ratio can be small. Another attempt on obtaining high-probability results for non-monotone functions could be to use some concentration analyses for submodular functions, such as [23]. However, the concentration analyses in prior works usually require the Lipschitz constant (the largest absolute marginal value) of the submodular function $f$ to be small compared to $f(O^{*})$. It remains open how to obtain high-probability results for maximizing general non-monotone submodular functions under noise.

### v2

@@ 4.2 Unconstrained Submodular Maximization
Finally, we consider maximizing any (non-monotone) submodular function without constraints. Here, we instantiate Theorem 3.4 with the double greedy algorithm [Buchbinder_Feldman_Seffi_Schwartz_2015] which is known to give a $1/2$-approximation for unconstrained submodular maximization. At a high-level, the double greedy algorithm works as follows. We first initialize two sets $X_{0}=\emptyset$ and $Y_{0}=N$. Let $N=\{e_{1},\ldots,e_{n}\}$. For $t=1,\ldots,n$, we check the marginal value $a_{t}$ of adding $e_{t}$ to $X_{t-1}$ and the marginal value $b_{t}$ of removing $e_{t}$ from $X_{t-1}$. With probability $a_{t}/(a_{t}+b_{t})$ (with some clipping operations if necessary), we add $e_{t}$ to $X_{t-1}$ to get $X_{t}$ and set $Y_{t}=Y_{t-1}$. Otherwise, we set $X_{t}=X_{t-1}$ and remove $e_{t}$ from $Y_{t-1}$ to get $Y_{t}$. A formal description of the algorithm can be found in Appendix D. The key fact that we require is the following lemma, with proof given in Appendix D.
@@
Fix $\varepsilon\in(0,\frac{2\nu^{2}}{\alpha})$. Suppose $n\geq\Omega(\frac{1}{\varepsilon}\log^{2}(\frac{n}{\varepsilon}))$. By letting the $\mathcal{A}$ in Algorithm 1 be the double greedy algorithm [Buchbinder_Feldman_Seffi_Schwartz_2015], we obtain a polynomial-time algorithm for maximizing unconstrained non-monotone submodular functions with noisy oracle satisfying $\mathbb{E}[f(\operatorname{ALG})]~\geq~\big(\frac{1}{2}-O(\varepsilon)\big)f(O^{*})$.
@@ 5 Simulation Results
  • Double greedy (DG) with exact value oracle [Buchbinder_Feldman_Seffi_Schwartz_2015]: this is a worst-case optimal polynomial-time algorithm in the noiseless setting. It uses the exact value oracle and is used only for reference.
@@ Appendix B Missing Proofs from Section 3
Lemma B.1 (Hoeffding’s inequality for sub-exponential distributions: see, e.g., page 29 in [wainwright_high-dimensional_2019]).
@@
Fix a set $S\subseteq N\setminus H$. Let $\mu=F^{H,t}(S)=\mathbb{E}_{H^{\prime}\sim H[t]}f(S\cup H^{\prime})=\frac{1}{\binom{h}{t}}\sum_{H^{\prime}\in H[t]}f(S\cup H^{\prime})$. By definition, $H_{1},\ldots,H_{m}$ are sampled from $H[t]$ without replacement. Since each $f(S\cup H^{\prime})$ is bounded in $0,f_{\max}]$, by Hoeffding’s inequality for sampling without replacement (see, e.g., Proposition 1.2 in [bardenet_concentration_2015]), we have
@@
11Use pipage rounding [calinescu_maximizing_2011] (which does not require access to $\hat{f}$) to convert the fractional solution $x(1)$ to a discrete set $S\in\mathcal{I}$.
@@
Algorithm 2 Measured continuous greedy [feldman_unified_2011] with approximate value oracle
In this section, we establish the robustness of the measured continuous greedy algorithm [feldman_unified_2011] (the full algorithm is given in Algorithm 2). Let $\hat{f}$ be an $\varepsilon$-approximate value oracle for $f$. The proof for the approximation ratios of the measured continuous greedy algorithm with exact value oracle is provided by FeldmanThesis2010. We analyze how the $\varepsilon$-approximate value oracle $\hat{f}$ will affect the approximation ratios. The main step is Lemma C.4 which provides a replacement of Corollary 3.2.7 of [FeldmanThesis2010]. Once this is established, we can apply the remaining arguments in Section 3.2.1 and Section 3.2.2 of [FeldmanThesis2010] by replacing the discretization error of $O(n^{3}\delta^{2})$ with the discretization and approximation error established in Lemma C.4.
@@ C.2 The Non-Monotone Case
The following lemma can be established from following the proofs of Lemma 3.2.8, Lemma 3.2.9, Corollary 3.2.10, and Lemma 3.2.11 from [FeldmanThesis2010] verbatim but replacing $O(n^{3}\delta)$ in their argument with $O(\alpha n+n^{3}\delta)$ (i.e. Lemma C.4).
@@ 7   With probability $\hat{p}_{i}$, let $X_{i}=X_{i-1}\cup\{u_{i}\},~Y_{i}=Y_{i-1}$;
Algorithm 3 Double greedy [Buchbinder_Feldman_Seffi_Schwartz_2015] with approximate value oracle
This section proves the robustness of the double greedy algorithm [Buchbinder_Feldman_Seffi_Schwartz_2015] (given in Algorithm 3) against approximate value oracle.
@@
In this case, it can be easily verified that $r_{i}=\max\{p_{i}b_{i},q_{i}a_{i}\}-\tfrac{1}{2}(p_{i}a_{i}+q_{i}b_{i})\leq 0$. Due to approximate oracle, $r_{i}=\max\{\hat{p}_{i}b_{i},\hat{q}_{i}a_{i}\}-\tfrac{1}{2}(\hat{p}_{i}a_{i}+\hat{q}_{i}b_{i})$ may be positive. [roughgarden_optimal_2018] show that the loss of performance of the double greedy algorithm due to approximate oracle can be upper bounded by $\sum_{i=1}^{n}r_{i}$:
Lemma D.2 (see Theorem 2.1 in [roughgarden_optimal_2018] or Lemma 4.3 in [Harvey_Liaw_Soma_2020]).
@@
    On the one hand, $b\leq\hat{b}+\varepsilon_{b}\leq\varepsilon_{b}$. On the other hand, because $a+b\geq 0$ holds for any submodular function [Buchbinder_Feldman_Seffi_Schwartz_2015], we have $a\geq-b\geq-\varepsilon_{b}$. Therefore,
@@
    On the one hand, $a\leq\hat{a}+\varepsilon_{a}\leq\varepsilon_{a}$. On the other hand, because $a+b\geq 0$ holds for any submodular function [Buchbinder_Feldman_Seffi_Schwartz_2015], we have $b\geq-a\geq-\varepsilon_{a}$. Therefore,
@@ Appendix F Discussion on High Probability Results
For monotone submodular functions, huang_efficient_2022 provide a method to do noisy comparison. For any set $S\subseteq N$, define the following comparison surrogate function $f_{0}(S)$ and the noisy version $\tilde{f}_{0}(S)$:
@@
For non-monotone submodular functions, the above approach to obtaining high-probability result does not work because: (1) the $f(S_{i})\geq f_{0}(S_{i})$ step in the noisy comparison analysis no longer holds, and (2) the size of the solution $|S_{i}|$ can be less than the rank $r$ of the matroid, so the $(1-\frac{1}{|S_{i}|})$ factor in the approximation ratio can be small. Another attempt on obtaining high-probability results for non-monotone functions could be to use some concentration analyses for submodular functions, such as [vondrak_note_2010]. However, the concentration analyses in prior works usually require the Lipschitz constant (the largest absolute marginal value) of the submodular function $f$ to be small compared to $f(O^{*})$. It remains open how to obtain high-probability results for maximizing general non-monotone submodular functions under noise.